A university graduate standing on the ground fires a rocket into the atmosphere at a velocity of 4000 m/s at an angle of 20o to the vertical axis. Determine (g = 9.8 m/s2):

(a) The time of flight (3) (b) The maximum level


1
Expert's answer
2022-03-25T09:11:21-0400

(a) the time of flight

"t=\frac{2v_0\sin\theta}{g}"

"t=\frac{2\times 4000\times \sin70^\circ}{9.8}=767\:\rm s"

(b) the maximum level

"H_{\max}=\frac{v_0^2\sin^2\theta}{2g}"

"H_{\max}=\frac{ 4000^2\times \sin^270^\circ}{2\times9.8}=7.2\times 10^5\:\rm m"